A carbon, oxygen and hydrogen containing compound has a composition of 69.6% C and 8.34% H by mass and has a molecular weight of 362.465 g/mol. What is its molecular formula of the compound?


1
Expert's answer
2021-10-26T05:35:44-0400

First get the percentage composition of oxygen.

= C + H + O = 100% 69.6 + O + 8.34 = 100

O= 100 - 77.9 = 22.06

Then calculate mass and moles of each

Carbon: 100% = 362.465 g/mol

therefore 69.6 will be = 69.6%/100% * 362.465 g/mol = 252.276 g/mol

divide by molar mass of carbon 252.276 g/mol/12g = 21


Hydrogen = 8.34/100 * 362.465 = 30.230/1 = 30


Oxygen = 22.06/100 * 362.465 = 79.960/16 = 5


formula = C21 H30 O5